# Global well-posedness of 2D Navier-Stokes with Dirichlet boundary fractional noise

### v1

@@ 1.2. Overview
The paper is organized as follows. In section 2 we introduce the functional setting to deal with problem (1.2). The proof of Theorem 1.3 is the object of section 3 and section 4. In particular, the global well-posedness, i.e. item 1, is addressed in section 3 considering the linear problem (1.3) in subsection 3.1 and the nonlinear one (1.4) in subsection 3.2. The interior regularity, i.e. item 2, is the object of section 4. In particular, in subsection 4.1 we study the interior regularity of the solution of the linear problem (1.3), while in subsection 4.2 we consider the nonlinear problem (1.4).
@@ 1.3. Notation
Here we collect some notation which will be used throughout the paper. Further notation will be introduced where needed. By $C$ we will denote several constants, perhaps changing value line by line. If we want to keep track of the dependence of $C$ from some parameter $\xi$ we will use the symbol $C(\xi)$. Sometimes we will use the notation $a\lesssim b$ (resp. $a\lesssim_{\xi}b$) if it exists a constant such that $a\leq Cb$ (resp. $a\leq C(\xi)b$).
Fix $q\in(1,\infty)$. For an integer $k\geq 1$, $W^{k,q}$ denotes the usual Sobolev spaces. In the non-integer case $s\in(0,\infty)\setminus{\mathbb{N}}$, we let $W^{s,q}=B^{s}_{q,q}$ where $B^{s}_{q,q}$ is the Besov space with smoothness $s$, and integrability $q$ and microscopic integrability $q$. Moreover, $H^{s,q}$ denotes the Bessel potential spaces. Both Besov and Bessel potential spaces can be defined using Littlewood-Paley decompositions and restrictions (see e.g. [56], [55, Section 6]) or using the interpolation methods starting with the standard Sobolev spaces $W^{k,q}$ (see e.g. [8, Chapter 6]). Finally, we set $\mathcal{A}^{s,q}(D;{\mathbb{R}}^{d})=(\mathcal{A}^{s,q}(D))^{d}$ for an integer $d\geq 1$, a domain $D$ and $\mathcal{A}\in\{W,H\}$. Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two Hilbert spaces. We denote by $\mathscr{L}_{2}(\mathcal{K}_{1},\mathcal{K}_{2})$ the set of Hilbert-Schmidt operators from $\mathcal{K}_{1}$ to $\mathcal{K}_{2}$. Below, we need the following Fubini-type result:
$H^{s}(D;\mathcal{K}_{1})=\mathscr{L}_{2}(\mathcal{K}_{1},H^{s}(D))\ \ \text{ for all }s\in{\mathbb{R}}.$
The above follows from [37, Theorem 9.3.6] and interpolation.
@@ 2.1. The Stokes operator and its spectral properties
We begin by introducing the Helmholtz projection on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$, see e.g. [53, Subsection 7.4]. Let $f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ and let $\psi_{f}\in W^{1,r}(\mathcal{O})$ be the unique solution of the following elliptic problem
(2.1) $\Delta\psi_{f}&=\operatorname{div}f\quad&\text{ on }&\mathcal{O},\\ \partial_{n}\psi_{f}&=f\cdot\widehat{n}&\text{ on }&\Gamma_{u}\cup\Gamma_{b}.\end{aligned}\right.$
Here $n$ denotes the exterior normal vector field on $\partial\mathcal{O}$. Of course, the above elliptic problem is interpreted in its naturally weak formulation:
@@
By [53, Corollary 7.4.4] , we have $\psi_{f}\in W^{1,r}(\mathcal{O})$ and $\|\nabla\psi_{f}\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}\lesssim\|f\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}$. Then the Helmholtz projection is given by $\mathbb{P}_{r}$ is defined as
@@
Next, we define the Stokes operator on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$. For notational convenience, we define $A_{r}$ as minus the Stokes operator so that $A_{r}$ is a positive operator for $r=2$ (i.e. $\langle A_{2}u,u\rangle\geq 0$ for all $u\in\mathsf{D}(A_{2})$). Let
@@
In the main arguments, we need stochastic and deterministic maximal $L^{r}$-regularity estimates for convolutions. By [46, 40], it is enough to provide the boundedness of the $H^{\infty}$-calculus for $A_{r}$. The reader is referred to [53, Chapters 3 and 4] and [37, Chapter 10] for the main notation and basic results on the $H^{\infty}$-calculus.
Contrary to [1], the boundary conditions we are interested in here are much more classical. Indeed, the Stokes operator with no-slip boundary conditions is well-studied. The reader is referred e.g. to [36, Section 2.8], [47], [32] and [38, Section 9] for the proof of this nowadays classical statement.
@@
The above implies that $-A_{r}$ generates an analytic semigroup on $\mathbb{L}^{r}$ which admits stochastic and deterministic maximal $L^{p}$-regularity for all $p\in(1,+\infty)$, see [53, Chapter 3-4] and [46]. We denote such semigroup by $S_{r}(t)$. We continue introducing some known facts about the “Sobolev tower” of spaces associated with the operator $A_{r}$. We denote by
@@
where $\sim$ denotes the completion. Indeed, since $0\in\rho(A_{r})$ by 2.1, we have that $f\mapsto\|A_{r}^{{\alpha}}f\|_{\mathbb{L}^{r}}$ is a norm for all $\alpha<0$. Since $(A_{r})^{*}=A_{r^{\prime}}$, it follows that (see e.g. [4, Chapter 5, Theorem 1.4.9])
@@
For convenience of notation, we simply write $A,\ S(t)$ in place of $A_{2}$ and $S(t)$. Moreover we define
@@ 2.2. The Dirichlet map
For all $g\in H^{-\frac{1}{2}}(\Gamma_{u})$ there exists a unique $(u,\pi)\in H\times H^{-1}(\mathcal{O})/\mathbb{R}$ very weak solution of (2.4). Moreover $(u,\pi)$ satisfy
@@
To begin, recall that $H^{s}(\Gamma_{u};U)=\mathscr{L}_{2}(U,H^{s}(\Gamma_{u}))$ for all $s\in{\mathbb{R}}$, see subsection 1.3. Hence, due to the ideal property of Hilbert-Schmidt operators, it is enough to consider the scalar case $U={\mathbb{R}}$.
@@
Hence, the claim now follows from the description of the fractional power of $A$ in 2.1. ∎
@@ 2.3. Deterministic Navier-Stokes equations
for all $\phi\in X_{1/2,A_{q}}$. Moreover, when $u\cdot\nabla v\in L^{r}\left(\mathcal{O};\mathbb{R}^{2}\right)$, it is explicitly given by
@@
The well-posedness of (2.8) in the sense of 2.4 is a well-known fact. Indeed the following theorem holds, see for instance [43, 59, 60].
@@
We end this section with the following lemma, which is the analogous of [29, Lemma 1.14] to the $L^{p}_{t}L^{q}_{x}$ framework.
@@
If $p\geq\frac{2q}{q-2}$, $u\in C([0,T];H)\cap L^{2}(0,T;V),\ v\in L^{p}(0,T;\mathbb{L}^{q})$, then
@@ 2.4. Stochastic convolutions with fractional noise
Let $U$ be a separable Hilbert space. A $U$-cylindrical fractional Brownian motion $(W^{\mathcal{H}}(t))_{t\geq 0}$ with Hurst index $H\in(0,1)$ is defined by the formal series
@@
For our aims in 3.1, we need the following results on the regularity of stochastic convolutions established in [22, Corollary 3.1].
@@
([22, Corollary 3.1]) Let $A$ be the generator of an analytic $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on a separable Hilbert space $U_{1}$, $\Phi\in\mathcal{L}(U,U_{1})$. Assume that
(2.14) $\|S(t)\Phi\|_{\mathcal{L}_{2}(U_{1},U)}\leq t^{-\gamma}\text{ for}\gamma<\mathcal{H}.$
Then the stochastic convolution $\int_{0}^{t}S(t-s)\Phi~{}\mathrm{d}W^{\mathcal{H}}(s)$ has $\mathbf{P}$-a.s. $\gamma_{1}$-Hölder continuous trajectories in $\mathsf{D}(A^{\gamma_{2}})$, for $0\leq\gamma_{1}+\gamma_{2}<\mathcal{H}$. If $\Phi\in\mathcal{L}_{2}(U_{1},U)$, then the assumption (2.14) is satisfied for $\gamma=0$.
@@ 3. Global well-posedness
Here we prove Theorem 1.3(1). This section is organized as follows. Firstly, in subsection 4.1 we first prove that the solution $w_{g}$ of the 2D Stokes equations with boundary noise (1.3) satisfies (1.5). Secondly, in subsection 3.2, we prove the existence of a $q$-solution to (1.2) by studying the auxiliary Navier-Stokes problem (1.4) for a given forcing term $w=w_{g}$ satisfying the regularity assumption as in (1.5) for a given $q$. Finally, in subsection 3.3, we prove the uniqueness of solutions to (1.2) therefore concluding the proof of Theorem 1.3(1). Recall that ($q$-)solutions of (1.2) are defined in 1.2.
@@ 3.1. Stokes equations
As discussed in subsection 1.1, we start by considering the linear problem (1.3). According to [20] and [21, Chapter 15], the mild solution $w_{g}$ of the former problem is formally given by
@@
Here $A$ is (minus) the Stokes operator with homogeneous boundary conditions (see Subsection 2.1).
@@
Note that, thanks to 2.3
$\mathcal{D}g\in\mathcal{L}_{2}(U,\mathsf{D}(A^{\frac{1}{4}-\frac{s+\varepsilon}{2}})).$
Hence, by 2.8, a.s.,
@@
The arbitrariness of $\varepsilon>0$ yields (3.2). To prove (3.3), note that, by 2.1,
@@
In the setting of 3.1, we have $2\mathcal{H}-\frac{3}{2}-s<\frac{1}{2}.$ Therefore, for all choices of $\mathcal{H}$ and $s$ in 1.1, it follows that
@@
Thus, (3.3) holds with $r<4$ and therefore $B(w_{g},w_{g})\not\in L^{2}(0,T;V^{*}).$ In particular, in the next subsection, we cannot avoid the use of $L^{p}$-setting in space, cf., the comments below 3.4.
@@
Previous results with white noise boundary conditions [10], [1] exploited stochastical maximal $L^{p}$ regularity techniques to study the linear part of the problem. Here is worth mentioning that we employed the more standard Hilbert value framework because it produces the sharpest result on the regularity of the stochastic convolution in terms of the Hurst parameter $\mathcal{H}$. Indeed, assuming just for simplicity the case $s=0$ and $g\in L^{p}(\Gamma_{u};U)$ for some $p\in[2,+\infty)$, then by 2.3, [17, Proposition 4.5] and arguing as above we have

### v2

@@ 1.2. Overview
In Section 2, we introduce the functional framework required to study problem (1.2). The proof of Theorem 1.3 is developed in Section 3 and Section 4. Specifically, global well-posedness, i.e. Item 1, is addressed in Section 3, where we first analyze the linear problem (1.4) in Section 3.1, followed by the nonlinear problem (1.5) in Section 3.2. Interior regularity, i.e. Item 2, is the focus of Section 4. In particular, Section 4.1 is devoted to the interior regularity of the solution to the linear problem (1.4), while Section 4.2 deals with the nonlinear problem (1.5).
@@ 1.3. Notation
Here we collect some notation which will be used throughout the paper. Additional notation will be introduced where needed. We use $C$ to denote a generic constant, which may vary from line to line. When it is important to emphasize the dependence of $C$ on a parameter $\xi$, we write $C(\xi)$. Moreover, we sometimes write $a\lesssim b$ (resp. $a\lesssim_{\xi}b$) to mean that there exists a constant $C$ (resp. $C(\xi)$) such that $a\leq Cb$ (resp. $a\leq C(\xi)b$).
Let $q\in(1,\infty)$ be fixed. For any integer $k\geq 1$, we denote by $W^{k,q}$ the standard Sobolev spaces. In the case of non-integer smoothness $s\in(0,\infty)\setminus\mathbb{N}$, we define $W^{s,q}=B^{s}_{q,q}$ where $B^{s}_{q,q}$ is the Besov space with smoothness $s$, and integrability $q$ and microscopic integrability $q$. We also denote by $H^{s,q}$ the Bessel potential spaces. Both Besov and Bessel potential spaces can be defined via Littlewood–Paley theory (see, e.g., [58], [57, Section 6]), or through interpolation methods based on the classical Sobolev spaces $W^{k,q}$ (see, e.g., [8, Chapter 6]). For a domain $D\subset{\mathbb{R}}^{n}$, integer $d\geq 1$, and $\mathcal{A}\in\{W,H\}$, we define the vector-valued spaces by $\mathcal{A}^{s,q}(D;{\mathbb{R}}^{d})=(\mathcal{A}^{s,q}(D))^{d}$.
Let $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ be two separable Hilbert spaces. We denote by $\mathscr{L}_{2}(\mathcal{K}_{1},\mathcal{K}_{2})$ the set of Hilbert-Schmidt operators from $\mathcal{K}_{1}$ to $\mathcal{K}_{2}$. We will use the following Fubini-type identity:
$H^{s}(D;\mathcal{K}_{1})=\mathscr{L}_{2}(\mathcal{K}_{1},H^{s}(D))\ \ \text{ for all }s\in{\mathbb{R}},$
which follows from [39, Theorem 9.3.6] and interpolation theory.
@@ 2.1. The Stokes operator and its spectral properties
We begin by introducing the Helmholtz projection on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$, see e.g. [55, Subsection 7.4]. As described also in [1, Section 2.1], the projection can be defined via an elliptic problem. We recall its construction here for the sake of completeness. Let $f\in L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ and let $\psi_{f}$ be the unique solution of the following elliptic problem
(2.1) $\Delta\psi_{f}&=\operatorname{div}f\quad&\text{ on }&\mathcal{O},\\ \partial_{\widehat{n}}\psi_{f}&=f\cdot\widehat{n}&\text{ on }&\Gamma_{u}\cup\Gamma_{b}.\end{aligned}\right.$
Here $\widehat{n}$ denotes the exterior normal vector field on $\partial\mathcal{O}$. This problem is understood in its standard weak formulation:
@@
By [55, Corollary 7.4.4] , we have $\psi_{f}\in W^{1,r}(\mathcal{O})$ and $\|\nabla\psi_{f}\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}\lesssim\|f\|_{L^{r}(\mathcal{O};{\mathbb{R}}^{2})}$. The Helmholtz projection $\mathbb{P}_{r}:L^{r}(\mathcal{O};{\mathbb{R}}^{2})\rightarrow L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ is then defined by
@@
Next, we define the Stokes operator on $L^{r}(\mathcal{O};{\mathbb{R}}^{2})$ corresponding to the boundary conditions considered in (1.2). For notational convenience, we define $A_{r}$ as minus the Stokes operator so that $A_{r}$ is a positive operator for $r=2$ (i.e. $\langle A_{2}u,u\rangle\geq 0$ for all $u\in\mathsf{D}(A_{2})$). Let
@@
In the main arguments, we need stochastic and deterministic maximal $L^{r}$-regularity estimates for convolutions. By [48, 42], it is enough to provide the boundedness of the $H^{\infty}$-calculus for $A_{r}$. The reader is referred to [55, Chapters 3 and 4] and [39, Chapter 10] for the main notation and basic results on the $H^{\infty}$-calculus.
Contrary to [1], the boundary conditions we are interested in here are much more classical. Indeed, the Stokes operator with no-slip boundary conditions is well-studied. The reader is referred e.g. to [38, Section 2.8], [49], [34] and [40, Section 9] for the proof of this nowadays classical statement.
@@
The above implies that $-A_{r}$ generates an analytic semigroup on $\mathbb{L}^{r}$ which admits stochastic and deterministic maximal $L^{p}$-regularity for all $p\in(1,+\infty)$, see [55, Chapter 3-4] and [48]. We denote such semigroup by $S_{r}(t)$. We continue introducing some known facts about the “Sobolev tower” of spaces associated with the operator $A_{r}$. We denote by
@@
where $\sim$ denotes the completion. Indeed, since $0\in\rho(A_{r})$ by Lemma 2.1, we have that $f\mapsto\|A_{r}^{{\alpha}}f\|_{\mathbb{L}^{r}}$ is a norm for all $\alpha<0$. Since $(A_{r})^{*}=A_{r^{\prime}}$, it follows that (see e.g. [4, Chapter 5, Theorem 1.4.9])
@@
For notational convenience, we will write $A,\ S(t)$ instead of $A_{2}$ and $S(t)$. Moreover we define
@@
For the convenience of the reader, we recall here the definition of the deterministic maximal $L^{p}$-regularity, since our techniques heavily rely on it.
Definition 2.2.
([55, Definition 3.5.1]) We let $\mathcal{X}$ stand for a Banach space and $\mathcal{A}$ a linear closed operator on $\mathcal{X}$ with domain $\mathsf{D}(\mathcal{A})$. We say that the inhomogeneous initial value problem on $L^{p}(0,T;\mathcal{X})$ given by
(2.4) $\dot{u}(t)+\mathcal{A}u(t)=f(t),~{}~{}u(0)=u_{0}$
admits maximal $L^{p}$-regularity, if for each $f\in L^{p}(0,T;\mathcal{X})$ and $u_{0}\in(\mathcal{X},\mathsf{D}(\mathcal{A}))_{1-1/p,p}$ there exists a unique $u\in W^{1,p}(0,T;\mathcal{X})\cap L^{p}(0,T;\mathsf{D}(\mathcal{A}))$ satisfying (2.4) a.e. in $(0,T)$.
@@ 2.2. The Dirichlet map
For all $g\in H^{-\frac{1}{2}}(\Gamma_{u})$ there exists a unique very weak solution $(u,\pi)\in H\times H^{-1}(\mathcal{O})/\mathbb{R}$ of (2.5). Moreover $(u,\pi)$ satisfy
@@
To begin, recall that $H^{s}(\Gamma_{u};U)=\mathscr{L}_{2}(U,H^{s}(\Gamma_{u}))$ for all $s\in{\mathbb{R}}$, see Section 1.3. Hence, due to the ideal property of Hilbert-Schmidt operators, it is enough to consider the scalar case $U={\mathbb{R}}$.
@@
Hence, the claim now follows from the description of the fractional power of $A$ in Lemma 2.1. ∎
@@ 2.3. Deterministic Navier-Stokes equations
for all $\phi\in X_{1/2,A_{q}}$. Moreover, if $u\cdot\nabla v\in L^{r}\left(\mathcal{O};\mathbb{R}^{2}\right)$ for some $r\in(1,\infty)$, it is explicitly given by
@@
The well-posedness of (2.9) in the sense of Definition 2.5 is a well-known fact. Indeed the following theorem holds, see for instance [45, 62, 63].
@@
We end this section with the following lemma, which generalizes [31, Lemma 1.14] to the $L^{p}_{t}L^{q}_{x}$ setting. In particular, [31, Lemma 1.14] corresponds to the specific case $p=q=4.$
@@
If $q>2$ and $p\geq\frac{2q}{q-2}$, $u\in C([0,T];H)\cap L^{2}(0,T;V),\ v\in L^{p}(0,T;\mathbb{L}^{q})$, then
@@ 2.4. Stochastic convolutions with fractional noise
Let $U$ be a separable Hilbert space. A $U$-cylindrical fractional Brownian motion $(W^{\mathcal{H}}(t))_{t\geq 0}$ with Hurst index $\mathcal{H}\in(0,1)$ is defined by the formal series
@@
For our aims in Proposition 3.1, we need the following results on the regularity of stochastic convolutions established in [23, Corollary 3.1] and [24, Proposition 11.6].
@@
([23, Corollary 3.1] and [24, Proposition 11.6]) Let $A$ be the generator of an analytic $C_{0}$-semigroup $(S(t))_{t\geq 0}$ on a separable Hilbert space $U_{1}$, $\Phi\in\mathscr{L}(U,U_{1})$. Assume that
(2.15) $\|S(t)\Phi\|_{\mathscr{L}_{2}(U,U_{1})}\leq t^{-\gamma}\text{ for}\gamma<\mathcal{H}.$
Then the stochastic convolution $\int_{0}^{t}S(t-s)\Phi~{}\mathrm{d}W^{\mathcal{H}}(s)$ has $\mathbf{P}$-a.s. $\gamma_{1}$-Hölder continuous trajectories in $\mathsf{D}(A^{\gamma_{2}})$, for $0\leq\gamma_{1}+\gamma_{2}<\mathcal{H}-{\gamma}$. If $\Phi\in\mathscr{L}_{2}(U,U_{1})$, then the assumption (2.15) is satisfied for $\gamma=0$.
@@ 3. Global well-posedness
Here we prove Theorem 1.3(1). This section is organized as follows. Firstly, in Section 3.1 we prove that the solution $w_{g}$ of the 2D Stokes equations with boundary noise (1.4) satisfies (1.6). Secondly, in Section 3.2, we prove the existence of a $q$-solution to (1.2) by studying the auxiliary Navier-Stokes problem (1.5) for a given forcing term $w=w_{g}$ satisfying the regularity assumption as in (1.6) for a given $q$. Finally, in Section 3.3, we prove the uniqueness of solutions to (1.2) therefore concluding the proof of Theorem 1.3(1). Recall that ($q$-)solutions of (1.2) are defined in Definition 1.2.
@@ 3.1. Stokes equations
As discussed in Section 1.1, we start by considering the linear problem (1.4). According to [21] and [22, Chapter 15], the mild solution $w_{g}$ of the former problem is formally given by
@@
Here $A$ is (minus) the Stokes operator with homogeneous boundary conditions as defined in Section 2.1.
@@
Note that, thanks to Corollary 2.4
$\mathcal{D}g\in\mathscr{L}_{2}(U,\mathsf{D}(A^{\frac{1}{4}-\frac{s+\varepsilon}{2}})).$
Hence, by Lemma 2.9, a.s.,
@@
The arbitrariness of $\varepsilon>0$ yields (3.2). To prove (3.3), note that, by Lemma 2.1,
@@
In the setting of Proposition 3.1, we have $2\mathcal{H}-\frac{3}{2}-s<\frac{1}{2}.$ Therefore, for all choices of $\mathcal{H}$ and $s$ in 1.1, it follows that
@@
Thus, (3.3) holds with $r<4$ and therefore $B(w_{g},w_{g})\not\in L^{2}(0,T;V^{*}).$ In particular, in the next subsection, we cannot avoid the use of $L^{p}$-setting in space, cf. the comments below 3.6.
@@
Previous results with white noise boundary conditions [1, 10] exploited stochastical maximal $L^{p}$ regularity techniques to study the linear part of the problem. Here is worth mentioning that we employed the more standard Hilbert value framework because it produces the sharpest result on the regularity of the stochastic convolution in terms of the Hurst parameter $\mathcal{H}$. Indeed, assuming just for simplicity the case $s=0$ and $g\in L^{p}(\Gamma_{u};U)$ for some $p\in[2,+\infty)$, then by Corollary 2.4, [18, Proposition 4.5] and arguing as above we have
@@
We end this subsection showing a lemma concerning the relation between the mild and the weak formulation of (1.4) as defined below.
Definition 3.4.
$w_{g}\in C(0,T;\mathbb{L}^{r})$
As above, $\widehat{n}$ denotes the exterior normal vector field on $\partial\mathcal{O}$. Since $g$ is time-independent, the last term in (3.4) can be rewritten as a stochastic integral as
Lemma 3.5.
Let 1.1 be satisfied. There exists a unique weak solution of (1.4) in the sense of Definition 3.4 and it is given by the formula (3.1).
Proof.
We split the proof into two steps.
Step 1: There exists a unique weak solution of (1.4) and it is necessarily given by the mild formula (3.1). Let $\psi\in C^{1}([0,T];\mathsf{D}(A))$. Arguing as in the first step of the proof of [31, Theorem 1.7], see also [46, Lemma 3], one can readily check that $w_{g}$ satisfies
for each $t\in[0,T],\ \mathbf{P}-a.s.$ The stochastic integral in the relation above is well-defined as a real-valued stochastic integral. Indeed, recalling that $(W^{\mathcal{H}}_{t})_{t\geq 0}$ is a $U$-cylindrical fractional Brownian motion we observe that $\langle g,\widehat{n}\cdot\nabla\psi(s)\rangle_{H^{-s}(\Gamma_{u}),H^{s}(\Gamma_{u})}$ is given by the linear operator on $U$
In conclusion, the stochastic integral in (3.1) is well-defined as a real-valued one (see [23, (2.16)]), since
Step 2: The mild formula (3.1) is a weak solution of (1.4) in the sense of Definition 3.4. We begin by noticing that $w_{g}$ has the required regularity due to Proposition 3.1. Let us test our mild formulation (3.1) against functions $\phi\in\mathsf{D}(A^{2})$. It holds, exploiting self-adjointness property of $S$ and $A$
The relation (3.1) is true. Indeed,